# Overcoming the Combinatorial Bottleneck in Symmetry-Driven Crystal Structure Prediction

### 04 Mar 2026 v2

@@ Abstract
Crystal structure prediction (CSP), which aims to predict the three-dimensional atomic arrangement of a crystal from its composition, is central to materials discovery and mechanistic understanding. However, given the composition and atomic counts in a unit cell, existing methods struggle with the NP-hard combinatorial challenge of rigorous symmetry enforcement or rely on retrieving known templates, which inherently limits both physical fidelity and the ability to discover genuinely new materials. To solve this, we propose a symmetry-driven generative framework. Our approach leverages large language models to encode chemical semantics and directly generate fine-grained Wyckoff patterns from atomic stoichiometry and counts, effectively circumventing the limitations inherent to database lookups. Crucially, to overcome the exponentially complex problem of combinatorial site assignments, we incorporate domain knowledge through an efficient, linear-complexity heuristic beam search algorithm that rigorously enforces algebraic consistency between site multiplicities and atomic stoichiometry and counts. By integrating this symmetry-consistent template into a diffusion backbone, our approach constrains the stochastic generative trajectory to a physically valid geometric manifold. This framework achieves state-of-the-art performance across stability, uniqueness, and novelty (SUN) benchmarks, alongside superior matching performance, thereby establishing a new paradigm for the rigorous exploration of targeted crystallographic space which can be previously uncharted, with no reliance on existing databases or a priori structural knowledge.
@@
However, rigorously enforcing these physical constraints imposed by space groups remains a universal bottleneck across the entire CSP domain. Fundamentally, determining an exact combination of Wyckoff positions whose multiplicities perfectly align with the exact atom counts of a given composition in the unit cell constitutes an inherently NP-hard combinatorial optimization problem, characterized by an exponentially exploding search space. Current methods struggle to efficiently resolve this task, often resorting to partial compromises. For example, aiming at efficiently addressing site assignments, Xin et al. [26] introduces a search method for Wyckoff positions with a pruning strategy; however, the approach remains a brute-force enumeration that inevitably faces this intractable exponential complexity for large atomic systems, and it lacks a likelihood-based ranking for the generated candidates.
@@
Figure 2: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., BaP3 ($N=16$). The definition for Stability, Uniqueness, and Novelty are detailed in Appendix C.1.
@@
Figure 3: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., K2AgBiI6 ($N=40$). The definition for Stability, Uniqueness, and Novelty are detailed in Appendix C.1.
(a) Retrieved and Prediction results of DiffCSP++.
(b) Prediction results of our method.
Figure 4: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., Lu2AlRu ($N=16$). The definition for Stability, Uniqueness, and Novelty are detailed in Appendix C.1.
@@
As shown in Table 1, our method yields notable and consistent improvements over DiffCSP++ across all three benchmarks on the comprehensive SUN metrics. On the MP-20 dataset, our method improves stability and novelty by approximately 124% and 71% respectively relative to the baseline. The advantages are even more pronounced on the challenging MPTS-52 dataset, where our method achieves a relative gain of about 255% in stability and 53% in novelty. Crucially, these component-level improvements compound to yield a striking 376% increase in the overall SUN metric. Even on Perov-5, which features relatively simple structural motifs, our method maintains a clear advantage in stability while matching the saturated uniqueness and novelty scores of the baseline. The significant rise in the overall SUN metric across all benchmark datasets confirms that our method generates candidates that are simultaneously more physically stable and structurally novel, while preserving the high standards of uniqueness. Figs. 2-4 visualize the prediction results for representative testing samples: BaP3, K2AgBiI6, and Lu2AlRu. The baseline DiffCSP++ forces these compositions into retrieved, topologically mismatched Wyckoff templates, inducing prominent atomic clashes. Consequently, these baseline structures fail both the stability and novelty criteria. In contrast, our approach infers compatible space groups and Wyckoff assignments directly from the stoichiometry. This resolves the structural conflicts and generates candidate materials that successfully satisfy all three criteria: Stability, Uniqueness, and Novelty.
@@
The performance disparity observed in Table  1 and Fig. 2-4 stems from the fundamental transition from retrieval-based memorization to generative reasoning. The baseline relies on CSPML, which essentially operates as a static lookup system. It rests on the isomorphism assumption, presupposing that for any new material, a structurally identical twin already exists in the training database. Consequently, this approach fails in data-sparse regimes or when exploring truly novel materials where no such prototype exists. When the model forces a new querying composition into a retrieved but topologically mismatched template, it introduces geometric frustration. The atoms are coerced into energetically unfavorable coordination environments to satisfy incorrect symmetry constraints, directly leading to poor stability. This reliance on data lookup inherently confines the model to the known structural motifs, rendering it incapable of discovering new polymorphs that lie outside the existing library.
@@
In stark contrast, our method reformulates symmetry determination as ab initio generative inference coupled with rigorous constrained optimizationInstead of retrieving existing solutions, the model learns a unified representation space from diverse structural data, enabling it to generalize and derive valid Wyckoff patterns for unseen compositions based on chemical rules. Crucially, a heuristic combinatorial optimization algorithm enforces algebraic consistency, ensuring the sum of site multiplicities matches the atomic stoichiometry. This mechanism empowers the model to explore uncharted chemical space while confining stochastic diffusion to a symmetry-compliant configurational space, guaranteeing output rationality. By eliminating symmetry-breaking artifacts of unconstrained generation, our framework enables inverse design of materials with high structural novelty and strict physical rigor.

### 07 Mar 2026 v3

# Symmetry-Driven Generation of Crystal Structures from Composition

@@ Abstract
Crystal structure prediction (CSP), which aims to predict the three-dimensional atomic arrangement of a crystal from its composition, is central to materials discovery and mechanistic understanding. However, given the composition in a unit cell, existing methods struggle with the NP-hard combinatorial challenge of rigorous symmetry enforcement or rely on retrieving known templates, which inherently limits both physical fidelity and the ability to discover genuinely new materials. To solve this, we propose a symmetry-driven generative framework. Our approach leverages large language models to encode chemical semantics and directly generate fine-grained Wyckoff patterns from atomic stoichiometry, effectively circumventing the limitations inherent to database lookups. Crucially, to overcome the exponentially complex problem of combinatorial site assignments, we incorporate domain knowledge through an efficient, linear-complexity heuristic beam search algorithm that rigorously enforces algebraic consistency between site multiplicities and atomic stoichiometry. By integrating this symmetry-consistent template into a diffusion backbone, our approach constrains the stochastic generative trajectory to a physically valid geometric manifold. This framework achieves state-of-the-art performance across stability, uniqueness, and novelty (SUN) benchmarks, alongside superior matching performance, thereby establishing a new paradigm for the rigorous exploration of targeted crystallographic space which can be previously uncharted, with no reliance on a priori structural knowledge.
@@
However, rigorously enforcing these physical constraints imposed by space groups remains a universal bottleneck across the entire CSP domain. Fundamentally, determining an exact combination of Wyckoff positions whose multiplicities perfectly align with the exact atom counts of a given composition in the unit cell constitutes an inherently NP-hard combinatorial optimization problem, characterized by an exponentially exploding search space. Current methods struggle to efficiently resolve this task, often resorting to partial compromises. For example, aiming at efficiently addressing site assignments, Xin et al. [26] introduces a search method for Wyckoff positions with a pruning strategy. However, the approach remains fundamentally a brute-force enumeration. As the number of atoms in the unit cell scales up, the possible integer partitions of these atoms into available Wyckoff orbits explode exponentially. Consequently, for large atomic systems containing dozens of atoms, this exhaustive search encounters a severe combinatorial explosion, rendering the enumeration computationally prohibitive. Furthermore, the approach lacks a likelihood-based ranking mechanism, making it practically impossible to identify the most physically plausible candidates from the massive, unguided pool of generated combinations.
@@
To generate the 20 candidates for each test composition, we employ a hierarchical sampling strategy to ensure both macroscopic and microscopic structural diversity. First, we select the Top-5 most probable space groups predicted by $LLM_{g}$. Subsequently, for each selected space group, we extract the Top-4 valid Wyckoff assignments using our constrained beam search algorithm. This combination yields exactly 20 distinct symmetry templates per composition, which are then used to guide the diffusion generation process.
@@
Figure 2: Absolute frequency distribution (raw counts) of space groups among the candidates across the testing set of MP-20. Counts are calculated using a unique-counting principle per composition, where any specific space group present in the candidate structures for a single testing composition is counted exactly once. Our generative approach yields a profoundly larger total unique count, demonstrating a consistently higher richness in symmetry exploration per composition compared to the localized redundancy of the retrieval baseline, i.e. CSPML [14].
@@
Figure 3: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., Zr3Fe8Mo ($N=36$). The definition for stability, uniqueness, and novelty are detailed in Appendix C.1.
@@
Figure 4: Visual comparison of prediction results from DiffCSP++ and our method for a testing sample, i.e., NaCr4O8 ($N=26$). The definition for stability, uniqueness, and novelty are detailed in Appendix C.1.
@@
As shown in Table 1, our method yields notable and consistent improvements over DiffCSP++ across all three benchmarks on the comprehensive SUN metrics. On the MP-20 dataset, our method improves stability and novelty by approximately 124% and 71% respectively relative to the baseline. The advantages are even more pronounced on the challenging MPTS-52 dataset, where our method achieves a relative gain of about 255% in stability and 53% in novelty. Crucially, these component-level improvements compound to yield a striking 376% increase in the overall SUN metric. Even on Perov-5, which features relatively simple structural motifs, our method maintains a clear advantage in stability while matching the saturated uniqueness and novelty scores of the baseline. The significant rise in the overall SUN metric across all benchmark datasets confirms that our method generates candidates that are simultaneously more physically stable and structurally novel, while preserving the high standards of uniqueness.
As case studies, Figs. 3 and 4 visualize the prediction results for two representative testing samples: Zr3Fe8Mo and NaCr4O8. These examples expose an inherent generative dilemma in the retrieval-based DiffCSP++ baseline: its predicted structures systematically fail either the novelty or stability criteria. For Zr3Fe8Mo, the baseline merely retrieves an exact match of a documented non-centrosymmetric $R3m$ phase, which guarantees stability but completely fails the novelty criterion. Furthermore, it forces the 36 atoms into a fragmented, lower-symmetry template comprising multiple $3a$ and $9b$ orbits. In contrast, our symmetry-driven approach bypasses empirical retrieval to discover a genuinely novel centrosymmetric phase under the higher-symmetry $R\bar{3}m$ space group by elegantly allocating the atoms to higher-multiplicity orbits such as $18h$ and $6c$. Conversely, the baseline’s prediction for NaCr4O8 is novel but highly unstable because it rigidly assigns the framework-forming Cr atoms to zero-degree-of-freedom (0-DOF) special sites ($2a$, $2b$, and $4e$) under the $C2/m$ space group. This mathematically locks the coordinates, preventing local relaxation and inducing severe microscopic strain within the Cr–O octahedral network. Our method elegantly resolves this by allocating the Cr atoms to distinct $4i$ orbits, whose two internal degrees of freedom ($x$ and $z$) provide the critical geometric flexibility to adaptively optimize bond lengths and navigate toward a thermodynamically stable local minimum. Ultimately, by inferring compatible space groups and Wyckoff assignments directly from the stoichiometry, our approach effectively overcomes these generative bottlenecks to yield candidate materials that successfully satisfy all three criteria: stability, uniqueness, and novelty.
@@
The performance disparity observed in Table  1 and Fig. 34 stem from the fundamental transition from retrieval-based memorization to generative reasoning. The baseline relies on CSPML, which essentially operates as a static lookup system. It rests on the isomorphism assumption, presupposing that for any new material, a structurally identical twin already exists in the training database. Consequently, this approach fails in data-sparse regimes or when exploring truly novel materials where no such prototype exists. When the model forces a new querying composition into a retrieved but topologically mismatched template, it introduces geometric frustration. The atoms are coerced into energetically unfavorable coordination environments to satisfy incorrect symmetry constraints, directly leading to poor stability. This reliance on data lookup inherently confines the model to the known structural motifs, rendering it incapable of discovering new polymorphs that lie outside the existing library.
@@
In stark contrast, our method reformulates symmetry determination as ab initio generative inference coupled with rigorous constrained optimization. Instead of retrieving existing solutions, the model learns a unified representation space from diverse structural data, enabling it to generalize and derive valid Wyckoff patterns for unseen compositions based on chemical rules. Crucially, a heuristic combinatorial optimization algorithm enforces algebraic consistency, ensuring the sum of site multiplicities matches the atomic stoichiometry. This mechanism empowers the model to explore uncharted chemical space while confining stochastic diffusion to a symmetry-compliant configurational space, guaranteeing output rationality. By eliminating symmetry-breaking artifacts of unconstrained generation, our framework enables inverse design of materials with high structural novelty and strict physical rigor.
To further quantify generative symmetry diversity, we evaluate the distribution of unique space group instances generated across the testing set, as shown in Fig. 2. To faithfully capture the true exploratory richness for each individual composition, we apply a strict unique-counting principle: a unique space group instance is defined such that if a specific space group appears within the candidate structures generated for a single testing composition, it is counted exactly once for that composition, regardless of its redundancy. The results demonstrate that the CSPML retrieval baseline [14] often collapses into retrieving multiple structural templates that share the identical space group for a given composition. Because of this localized redundancy, multiple retrieved candidates frequently contribute only a single unique space group count, which drastically shrinks its total aggregate score. In contrast, our generative approach explicitly infers multiple mutually distinct space groups directly from the probability distribution predicted by the language models. This mechanism yields a profoundly larger total count of unique instances and demonstrates a consistently richer and more abundant exploration across different space groups for each of the composition, as evidenced by the higher frequency observed across nearly all space group indices in Fig. 2.